Detia could row 12 miles downstream in one hour less than she could row 6 miles upstream what was the speed of the current in the river and how fast could detia row in still water if the speed of the current was one half the speed detia could row in still water

Answer by mananth(16946)

Detia could row 12 miles downstream in one hour less than she could row 6 miles upstream what was the speed of the current in the river and how fast could detia row in still water if the speed of the current was one half the speed detia could row in still water


downstream distance = 12 miles
time taken = t-1
speed of boat = 2x
speed of current = x
downstream speed = 2x+x=3x

Upstream distance = 6 miles
Time taken = t
Upstream speed = 2x-x=x
Time upstream - time downstream = 1
6/x - 12/3x= 1
(18-12)/3x =1
6=3x
x=2
current speed = 2 mph
boat speed = 4 mph